RC Meeting and Runplan, Wednesday, Sept 26

From HallCWiki
Jump to: navigation, search
  • Status
    • Finished the SHMS beam deflection test during owl. Ion chamber calibrations good up to 20 uA. Beam established up to 50 uA.
    • Steve Lassiter fixing HMS Q2 this morning.
    • Target cooled and condensed this morning.
    • Day shift experts minimizing beam spot today and improving high current operations.
  • Runplan
1. Get beam into Hall with good beam spot size, push to high current with experts on day. 
2. Check beam spot with harps, beam on yag with and without raster, carbon hole (5 uA) 
3. Trigger and detector checkout [1] (can be done up to 20 uA--we have a current target ion chamber calibration good up to this current). 
4. New ion chamber calibrations up to highest current settings with target for optics and future running.
5. HMS/SHMS optics study (under Calibration Runs on pg 4 of run plan). 
  • Accesses
    • Steve Lassiter is going to take a look at the SHMS encoder tomorrow when beam is down for pass change.
  • Notices
    • Encoder angles in the rotation GUI drift, and users can no longer modify the offsets themselves.
    • JP will do target training 4pm Thursday. Contact him if interested!!
    • Hall A is planning a pass change for Thursday- there will be a few hours of no beam during that time.